You are planning to buy a good quality cell phone in order to attend the online calss.The average price of 50 cellphones is ₱13,500 with a standard deviation of ₱573.65 and a confidence level of 99%. Find the corresponding confidence interval.


1
Expert's answer
2022-05-30T03:43:10-0400

The critical value for "\alpha = 0.01," and "df = n-1 = 49" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.679952."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(13500-2.679952\times\dfrac{573.65}{\sqrt{50}},"

"13500+2.679952\times\dfrac{573.65}{\sqrt{50}})"

"=(13282.585, 13717.415)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "13282.585 < \mu < 13717.415," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(13282.585, 13717.415)."